Find: (2/5) ÷ 1 whole ½ Find: (2/5) ÷ 1 whole ½ Find: (2/5) ÷ 1 whole ½

Find:
 (2/5) ÷ 1 whole ½ 
Find:
 (2/5) ÷ 1 whole ½ 
Find:
 (2/5) ÷ 1 whole ½ 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

First covert the mixed fraction into improper fraction.
We have,
= 1 whole 1⁄2 = 3/2
Then,
= (2/5) × reciprocal of (3/2)
= (2/5) × (2/3)
= (2 × 2) / (5 × 3)
= 4/15

Thanks